OpenStudy (anonymous):

Find, if any, the local maximum, local minimum, and inflection points for y’ = 2x^2 – x – 3.

OpenStudy (amistre64):

y' is correct right? (2x-3)(x+1) x = 3/2 and x=-1 are bending points y'' = 4x-1 at x=-1 we have a (-)value its a MAX at x = 3/2 we will have a (+) value..its a MIN at x = 1/4 we should have an inflection